OpenStudy (anonymous):

Help Please ! Explain The Steps .. Simplify the sum. State any restrictions on the variables. (x - 2)/(x + 3) + (10x)/(x^2 - 9)

OpenStudy (anonymous):

\[\frac{ x - 2 }{ x + 3 } + \frac{ 10x }{ x^2 - 9 }\]

OpenStudy (anonymous):

@Hero please help

OpenStudy (anonymous):

@campbell_st help !!

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

walk me through this please

OpenStudy (campbell_st):

ok... so you need a common denominator to add fractions if you factorise the denominator of the 2nd fraction you get \[\frac{x -2}{x -3} + \frac{10x}{(x - 3)(x +3)}\] the common denominator is (x -3)(x + 3) so multiply the numerator and deniminator of the 1st fraction byt (x + 3)\[\frac{(x -2) \times (x + 3)}{(x - 3)\times(x + 3)} + \frac{10x}{(x -3)(x + 3)}\] if you distribute and collect like terms in the numerator then factorise if possible you'll find there is a common factor that cancels.